# Evaluating LLMs for Demographic-Targeted Social Bias Detection: A Comprehensive Benchmark Study

### v1

@@ Abstract
Large-scale web-scraped text corpora used to train general-purpose AI models often contain harmful demographic-targeted social biases, creating a regulatory need for data auditing and developing scalable bias-detection methods. Although prior work has investigated biases in text datasets and related detection methods, these studies remain narrow in scope. They typically focus on a single content type (e.g., hate speech), cover limited demographic axes, overlook biases affecting multiple demographics simultaneously, and analyze limited techniques. Consequently, practitioners lack a holistic understanding of the strengths and limitations of recent large language models (LLMs) for automated bias detection. In this study, we present a comprehensive evaluation framework aimed at English texts to assess the ability of LLMs in detecting demographic-targeted social biases. To align with regulatory requirements, we frame bias detection as a multi-label task using a demographic-focused taxonomy. We then conduct a systematic evaluation with models across scales and techniques, including prompting, in-context learning, and fine-tuning. Using twelve datasets spanning diverse content types and demographics, our study demonstrates the promise of fine-tuned smaller models for scalable detection. However, our analyses also expose persistent gaps across demographic axes and multi-demographic targeted biases, underscoring the need for more effective and scalable auditing frameworks.
@@
Keywords: Social bias, Bias detection, Prompting, Fine-tuning
[…]
Evaluating LLMs for Demographic-Targeted Social Bias Detection: A Comprehensive Benchmark Study
[…]
1 MPI-SWS and Saarland University, Saarbrücken, Germany
2 Huawei Technologies France S.A.S.U., Paris, France
3 Huawei Technologies Company Ltd., Shenzhen, China
ayanm@mpi-sws.org, {chenfeihao, jinghui.li, jasmine.xwang}@huawei.com
Abstract content
1.  Introduction
@@
Large-scale web-scraped text corpora have significantly driven recent advances in general-purpose AI (GPAI) models. These corpora, however, are often minimally curated and can contain harmful social biases, e.g., hateful, toxic, or stereotypical content targeting various demographic axes Navigli et al. (2023). Consequently, training on such content can lead to unsafe model outputs and real-world harms Dodge et al. (2021); Vashney (2022).
@@
Detecting demographic-targeted social biases in data has become both a regulatory and technical requirement Tabassi (2023); European Union (2024). For example, the EU’s GPAI Code of Practice highlights the importance of documenting demographic-centric harms in training corpora. Effective data-level mitigation strategies Gallegos et al. (2024) also depend on reliably identifying biased instances. While prior studies have examined biases in text corpora Kreutzer et al. (2022); Luccioni and Viviano (2021); Dodge et al. (2021), they typically rely either on small-scale human annotations or on simplistic automated techniques. Manual inspection, however, poses psychological risks Steiger et al. (2021), and despite advances in LLMs, few works offer a comprehensive analysis of the state of the art in social bias detection.
@@
Compared to systematic benchmarks that examine LLMs for biased generation Parrish et al. (2022); Sun et al. (2024), detailed studies of their potential as tools for detecting social biases in text remain underexplored. Existing work is often limited across one or more dimensions: focusing on narrow demographics Wang et al. (2024), restricting attention to a single type of content such as hate Mathew et al. (2021), analyzing only specific contexts like certain subreddits Kumar et al. (2024), or analyzing only few methods such as zero-shot prompting Sun et al. (2024). Moreover, most studies overlook biases that target multiple demographic axes simultaneously, e.g., intersectionality. These gaps motivate a structured and comprehensive study of the current capabilities of recent LLMs in detecting demographic-targeted social biases.
@@
2.  Related work
Bias in LLMs. Several works have evaluated biases in LLMs, independently analyzing content types like stereotypes Nadeem et al. (2021); Parrish et al. (2022) and hate/toxic content Gehman et al. (2020). Recently, Li et al. (2023) also studied the fairness of ChatGPT in binary decision-making. Several benchmarks also analyzed stereotype and toxic characteristics in generations of recently developed LLMs Wang et al. (2023); Sun et al. (2024); Wang et al. (2024).
Bias detection with LLMs. Prior work explored using LLMs in hate-speech moderation Barikeri et al. (2021); Mathew et al. (2021); Kumar et al. (2024); Zhan et al. (2025) or domain-specific bias detection Raza et al. (2024). Recent work Sun et al. (2024); Wang et al. (2024) also benchmarked prompt-based bias detection. However, they are restricted to zero-shot settings, cover limited demographics and data. In contrast, we systematically evaluate across multiple LLM-based methods, a broader range of demographics, and holistically analyze across different content types.
Bias analysis of corpora. Other work has directly analyzed large text corpora. Kreutzer et al. (2022) employed human surveys on a small web-crawled subset to assess multilingual quality and offensive content. Lexicon-based approaches have been used to detect opinion biases in Wikipedia Hube and Fetahu (2018). Luccioni and Viviano (2021) subsampled Common Crawl to study sexual and hateful content using n-grams, BERT, and logistic regression, while Dodge et al. (2021) analyzed C4, linking sentiment toward racial groups to biased QA outcomes. Although these studies provide valuable insights, they rely on relatively simplistic detection methods. In contrast, our work offers a systematic evaluation of state-of-the-art LLMs for social bias detection, providing deeper insights that can complement and extend prior analyses of large-scale corpora.
LLM guardrails. LLMs have also been explored as guardrails for GPAI systems Markov et al. (2023); Inan et al. (2023); Chen et al. (2024a); Zeng et al. (2024), primarily to mitigate harmful user prompts and model-generated outputs. While effective for moderating AI systems, these models are not designed for systematically identifying biases in raw text. As we later show, they fail to capture subtle social biases in texts, highlighting the need for dedicated evaluations and methods.
3.  Framework setup
@@
BBQ Parrish et al. (2022) Stereo 7843
BEC-Pro Bartl et al. (2020) Occup 2580
CrowS-pairs Nangia et al. (2020) Stereo 1278
HateXplain Mathew et al. (2021) Hate/Tox 5437
ImplicitHate ElSherief et al. (2021) Hate/Tox 2876
RedditBias Barikeri et al. (2021) Stereo 9087
SBIC Sap et al. (2020) Hate/Tox 5243
StereoSet Nadeem et al. (2021) Stereo 1417
ToxiGen Hartvigsen et al. (2022) Hate/Tox 2096
WinoBias Zhao et al. (2018) Occup 3168
Winogender Rudinger et al. (2018) Occup 240
Winoqueer Felkner et al. (2023) Stereo 5516
@@
Unbiased UNB its always a mistaken identity because they think every black person is a criminal.Barikeri et al. (2021)
Single-axis RAC what are you talking about arabs are rich af. they got arab moneyBarikeri et al. (2021)
Multi-axis GEN,RAC I like my women like I like my black people DeadSap et al. (2020)
GEN,RAC interracism is a form of race war. it is black males waging sexual war of conquest and colonizing white wombs to make brown babiesElSherief et al. (2021)
@@
3.1.  Demographic-targeted taxonomy
Prior evaluation studies are limited in scope: (i) they narrowly focus on specific content types (e.g., hate Mathew et al. (2021)), (ii) adopt narrow context-specific taxonomies (e.g., subreddits Kumar et al. (2024)), or (iii) consider only a small set of demographic axes Wang et al. (2024). Hence, we lack a holistic view of the current capabilities for detecting social biases Navigli et al. (2023), i.e., “prejudices, stereotypes, and discriminatory attitudes against certain groups” embedded in texts.
To address the limitations, our framework employs a demographic-centered taxonomy that identifies the demographic axes targeted by biased texts. This approach aligns directly with anti-discrimination regulations and governance measures European Commission (2025), while also enabling the study of multi-axis biases: cases where texts simultaneously target multiple groups, an aspect often overlooked. Concretely, our taxonomy spans nine axes with differing legal recognition:
@@ 3.1.  Demographic-targeted taxonomy
    Broad recognition: Gender identity (GEN), Sexual orientation (SO), Disability (DIS), Age (AGE), Race/ethnicity (RAC), Nationality (NAT), and Religion (REL), all widely protected Congress (1964); EU FRA (2018).
@@
    Narrow recognition: Socioeconomic status (SES) and Physical appearance (PHY), protected in certain regional frameworks Viprey (2002); Klose et al. (2025).
@@
Any text not targeting one or more of these axes is considered “unbiased” (UNB) within our taxonomy. Each demographic axis serves as a prediction category, making the detection task twofold: (i) identify whether a text expresses demographic-targeted bias, and (ii) determine which demographics are targeted. Unlike prior benchmarks that treat bias detection as single-label or multi-class classification Mathew et al. (2021); Wang et al. (2024), our formulation supports multi-label prediction, capturing both single-axis (e.g., race only) and multi-axis (e.g., gender+race) biases (Table 2).
@@
Our taxonomy also considers a broad range of content types encoding social biases through implicit harms Blodgett et al. (2020), including: i) Stereotype descriptions that stereotype, misrepresent, or disparage identities, ii) Occupation–gender associations that stereotype, erase, or exclude gender identities, and Hate or toxic content targeting demographics through toxicity, derogation, or dehumanization. However, by centering on demographic axes rather than surface content types, we enable systematic characterization of which demographics are harmed, allow for analysis of multi-axis cases, and align directly with anti-discrimination and AI governance frameworks European Union (2024).
3.2.  Adapting existing datasets
@@
To enable comprehensive empirical evaluation in realistic settings, we adapt existing English datasets to align with our demographic-targeted social bias taxonomy. We surveyed widely used NLP datasets Gallegos et al. (2024), prioritizing diversity across demographic axes and harm types. Unlike prior benchmarks Wang et al. (2024), which often rely on fully GPT-generated categories (e.g., toxic text), we minimize synthetic data to reduce evaluation artifacts Koo et al. (2024); Maheshwari et al. (2024). [^1] The only exception is ToxiGen, although it has human-annotated GPT text, unlike Wang et al. (2024). Based on this review, we curated samples from twelve distinct datasets (Table 1).
@@
Since some datasets were originally designed for evaluating generative biases, we applied minor adaptations similar to Wang et al. (2024). For StereoSet, we concatenate the context with the stereotype options. For BBQ, we construct text instances by pairing disambiguated contexts with answers. For CrowS-Pairs, we retain only the “more biased” sentences. For SBIC, we derive a single label using the majority vote across annotators. For ToxiGen, we label texts as biased only when human annotator scores indicate bias.
@@
We also observed inconsistencies in how demographic groups are labeled across datasets, so we standardized labels for consistency. For instance, bias against “Arabs” or “Middle Eastern” identities, labeled inconsistently as RAC Nadeem et al. (2021) or REL Barikeri et al. (2021), is mapped to RAC, reserving REL for explicit religious references. Biases against national identities such as “Chinese” or “Mexican” are assigned to NAT. Bias against “Jewish” identity is annotated as both RAC and REL to reflect its ethnoreligious nature Litt (1961). We also improve regulatory alignment by disambiguating GEN and SO (e.g., transgender bias labeled as GEN) and denoting pregnancy-targeted bias under GEN rather than PHY Parrish et al. (2022). Labels outside our taxonomy (e.g., victim Sap et al. (2020)) are excluded.
@@
The resulting dataset contains 46,781 entries, substantially larger than comparable benchmarks (e.g., 11,004 samples in Wang et al. (2024)). Biased instances are more prevalent (around 70%), with most targeting a single demographic axis and roughly 12% of biased instances targeting multiple axes simultaneously. Among demographic targets, GEN, RAC, SO, and REL are most common, while PHY is least prevalent. Multi-axis biases most frequently combine {GEN, SO} or {GEN, RAC}.
@@
3.3.1.  Prompting
Brown et al. (2020) demonstrated that instruction-tuned LLMs can effectively perform a variety of tasks through textual prompting in zero-shot scenarios. Following Inan et al. (2023); Palla et al. (2025), our evaluation framework employs policy-based prompting for bias detection. Specifically, the prompt includes a policy detailing the bias detection task and our demographic-based social bias taxonomy. To enhance the bias detection capabilities within this prompting framework, we also assess the benefits of incorporating few-shot examples for in-context learning. Specifically, we utilize the retrieval-augmented (RAG) framework Chen et al. (2024a), where an embedding model identifies the most relevant in-context examples for each input instance from the training/development set.
Models. We consider several instruction-tuned models ranging from 8B to 72B parameters, e.g., GLM-4 GLM et al. (2024), Llama-3.1 Dubey et al. (2024), and Qwen-2.5 Yang et al. (2024). We also analyze the guardrail model Llama Guard-3 Inan et al. (2023) to explore if such models could directly be applied for general text bias detection. To perform RAG-based in-context sample selection, we use the BGE-M3 Chen et al. (2024b) model.
3.3.2.  Fine-tuning

### v2

@@ Abstract
Large-scale web-scraped text corpora used to train general-purpose AI (GPAI) models often contain harmful demographic-targeted social biases, creating a regulatory need for data auditing and developing scalable bias-detection methods. Although prior work has investigated biases in text datasets and related detection methods, these studies remain narrow in scope. They typically focus on a single content type (e.g., hate speech), cover limited demographic axes, overlook biases affecting multiple demographics simultaneously, and analyze limited techniques. Consequently, practitioners lack a holistic understanding of the strengths and limitations of recent large language models (LLMs) for automated and scalable bias detection in datasets. In this study, we present a comprehensive evaluation framework aimed at English texts to assess the ability of LLMs in detecting demographic-targeted social biases. To align with regulatory requirements, we frame bias detection as a multi-label task using a demographic-focused taxonomy. We then conduct a systematic evaluation with models across scales and techniques, including prompting, in-context learning, and fine-tuning. Using twelve datasets spanning diverse content types and demographics, our study demonstrates the promise of fine-tuned smaller models for scalable detection. However, our analyses also expose persistent gaps across demographic axes and multi-demographic targeted biases, underscoring the need for more effective and scalable auditing frameworks.
@@
Keywords Social bias $\cdot$ Bias detection $\cdot$ Prompting $\cdot$ Fine-tuning
1 Introduction
@@
Large-scale web-scraped text corpora have significantly driven recent advances in general-purpose AI (GPAI) models. These corpora, however, are often minimally curated and can contain harmful social biases, e.g., hateful, toxic, or stereotypical content targeting various demographic axes (Navigli et al., 2023). Consequently, training on such content can lead to unsafe model outputs and real-world harms (Dodge et al., 2021; Vashney, 2022).
@@
Detecting demographic-targeted social biases in data has become both a regulatory and technical requirement (Tabassi, 2023; European Union, 2024). For example, the EU’s GPAI Code of Practice highlights the importance of documenting demographic-centric harms in training corpora. Effective data-level mitigation strategies (Gallegos et al., 2024) also depend on reliably identifying biased instances. While prior studies have examined biases in text corpora (Kreutzer et al., 2022; Luccioni and Viviano, 2021; Dodge et al., 2021), they typically rely either on small-scale human annotations or on simplistic automated techniques. Manual inspection, however, poses psychological risks (Steiger et al., 2021), and despite advances in LLMs, few works offer a comprehensive analysis of the state of the art in social bias detection.
@@
Compared to systematic benchmarks that examine LLMs for biased generation (Parrish et al., 2022; Sun et al., 2024), detailed studies of their potential as tools for detecting social biases in text remain underexplored. Existing work is often limited across one or more dimensions: focusing on narrow demographics (Wang et al., 2024), restricting attention to a single type of content such as hate (Mathew et al., 2021), analyzing only specific contexts like certain subreddits (Kumar et al., 2024), or analyzing only few methods such as zero-shot prompting (Sun et al., 2024). Moreover, most studies overlook biases that target multiple demographic axes simultaneously, e.g., intersectionality. These gaps motivate a structured and comprehensive study of the current capabilities of recent LLMs in detecting demographic-targeted social biases.
@@
2 Related work
Bias in LLMs. Several works have evaluated biases in LLMs, independently analyzing content types like stereotypes (Nadeem et al., 2021; Parrish et al., 2022) and hate/toxic content (Gehman et al., 2020). Recently, Li et al. (2023) also studied the fairness of ChatGPT in binary decision-making. Several benchmarks also analyzed stereotype and toxic characteristics in generations of recently developed LLMs (Wang et al., 2023; Sun et al., 2024; Wang et al., 2024).
Bias detection with LLMs. Prior work explored using LLMs in hate-speech moderation (Barikeri et al., 2021; Mathew et al., 2021; Kumar et al., 2024; Zhan et al., 2025) or domain-specific bias detection (Raza et al., 2024). Recent work (Sun et al., 2024; Wang et al., 2024) also benchmarked prompt-based bias detection. However, they are restricted to zero-shot settings, cover limited demographics and data. In contrast, we systematically evaluate across multiple LLM-based methods, a broader range of demographics, and holistically analyze across different content types.
Bias analysis of corpora. Other work has directly analyzed large text corpora. Kreutzer et al. (2022) employed human surveys on a small web-crawled subset to assess multilingual quality and offensive content. Lexicon-based approaches have been used to detect opinion biases in Wikipedia (Hube and Fetahu, 2018). Luccioni and Viviano (2021) subsampled Common Crawl to study sexual and hateful content using n-grams, BERT, and logistic regression, while Dodge et al. (2021) analyzed C4, linking sentiment toward racial groups to biased QA outcomes. Although these studies provide valuable insights, they rely on relatively simplistic detection methods. In contrast, our work offers a systematic evaluation of state-of-the-art LLMs for social bias detection, providing deeper insights that can complement and extend prior analyses of large-scale corpora.
LLM guardrails. LLMs have also been explored as guardrails for GPAI systems (Markov et al., 2023; Inan et al., 2023; Chen et al., 2024a; Zeng et al., 2024), primarily to mitigate harmful user prompts and model-generated outputs. While effective for moderating AI systems, these models are not designed for systematically identifying biases in raw text. As we later show, they fail to capture subtle social biases in texts, highlighting the need for dedicated evaluations and methods.
3 Framework setup
@@
BBQ (Parrish et al., 2022) Stereo 7843
BEC-Pro (Bartl et al., 2020) Occup 2580
CrowS-pairs (Nangia et al., 2020) Stereo 1278
HateXplain (Mathew et al., 2021) Hate/Tox 5437
ImplicitHate (ElSherief et al., 2021) Hate/Tox 2876
RedditBias (Barikeri et al., 2021) Stereo 9087
SBIC (Sap et al., 2020) Hate/Tox 5243
StereoSet (Nadeem et al., 2021) Stereo 1417
ToxiGen (Hartvigsen et al., 2022) Hate/Tox 2096
WinoBias (Zhao et al., 2018) Occup 3168
Winogender (Rudinger et al., 2018) Occup 240
Winoqueer (Felkner et al., 2023) Stereo 5516
@@
Unbiased UNB its always a mistaken identity because they think every black person is a criminal.(Barikeri et al., 2021)
Single-axis RAC what are you talking about arabs are rich af. they got arab money(Barikeri et al., 2021)
Multi-axis GEN,RAC I like my women like I like my black people Dead(Sap et al., 2020)
GEN,RAC interracism is a form of race war. it is black males waging sexual war of conquest and colonizing white wombs to make brown babies(ElSherief et al., 2021)
@@
3.1 Demographic-targeted taxonomy
Prior evaluation studies are limited in scope: (i) they narrowly focus on specific content types (e.g., hate (Mathew et al., 2021)), (ii) adopt narrow context-specific taxonomies (e.g., subreddits (Kumar et al., 2024)), or (iii) consider only a small set of demographic axes (Wang et al., 2024). Hence, we lack a holistic view of the current capabilities for detecting social biases (Navigli et al., 2023), i.e., “prejudices, stereotypes, and discriminatory attitudes against certain groups” embedded in texts.
To address the limitations, our framework employs a demographic-centered taxonomy that identifies the demographic axes targeted by biased texts. This approach aligns directly with anti-discrimination regulations and governance measures (European Commission, 2025), while also enabling the study of multi-axis biases: cases where texts simultaneously target multiple groups, an aspect often overlooked. Concretely, our taxonomy spans nine axes with differing legal recognition:
@@ 3.1 Demographic-targeted taxonomy
    Broad recognition: Gender identity (GEN), Sexual orientation (SO), Disability (DIS), Age (AGE), Race/ethnicity (RAC), Nationality (NAT), and Religion (REL), all widely protected (Congress, 1964; EU FRA, 2018).
@@
    Narrow recognition: Socioeconomic status (SES) and Physical appearance (PHY), protected in certain regional frameworks (Viprey, 2002; Klose et al., 2025).
@@
Any text not targeting one or more of these axes is considered “unbiased” (UNB) within our taxonomy. Each demographic axis serves as a prediction category, making the detection task twofold: (i) identify whether a text expresses demographic-targeted bias, and (ii) determine which demographics are targeted. Unlike prior benchmarks that treat bias detection as single-label or multi-class classification (Mathew et al., 2021; Wang et al., 2024), our formulation supports multi-label prediction, capturing both single-axis (e.g., race only) and multi-axis (e.g., gender+race) biases (Table 2).
@@
Our taxonomy also considers a broad range of content types encoding social biases through implicit harms (Blodgett et al., 2020), including: i) Stereotype descriptions that stereotype, misrepresent, or disparage identities, ii) Occupation–gender associations that stereotype, erase, or exclude gender identities, and Hate or toxic content targeting demographics through toxicity, derogation, or dehumanization. However, by centering on demographic axes rather than surface content types, we enable systematic characterization of which demographics are harmed, allow for analysis of multi-axis cases, and align directly with anti-discrimination and AI governance frameworks (European Union, 2024).
3.2 Adapting existing datasets
@@
To enable comprehensive empirical evaluation in realistic settings, we adapt existing English datasets to align with our demographic-targeted social bias taxonomy. We surveyed widely used NLP datasets (Gallegos et al., 2024), prioritizing diversity across demographic axes and harm types. Unlike prior benchmarks (Wang et al., 2024), which often rely on fully GPT-generated categories (e.g., toxic text), we minimize synthetic data to reduce evaluation artifacts (Koo et al., 2024; Maheshwari et al., 2024). [^1] The only exception is ToxiGen, although it has human-annotated GPT text, unlike (Wang et al., 2024). Based on this review, we curated samples from twelve distinct datasets (Table 1).
@@
Since some datasets were originally designed for evaluating generative biases, we applied minor adaptations similar to (Wang et al., 2024). For StereoSet, we concatenate the context with the stereotype options. For BBQ, we construct text instances by pairing disambiguated contexts with answers. For CrowS-Pairs, we retain only the “more biased” sentences. For SBIC, we derive a single label using the majority vote across annotators. For ToxiGen, we label texts as biased only when human annotator scores indicate bias.
@@
We also observed inconsistencies in how demographic groups are labeled across datasets, so we standardized labels for consistency. For instance, bias against “Arabs” or “Middle Eastern” identities, labeled inconsistently as RAC (Nadeem et al., 2021) or REL (Barikeri et al., 2021), is mapped to RAC, reserving REL for explicit religious references. Biases against national identities such as “Chinese” or “Mexican” are assigned to NAT. Bias against “Jewish” identity is annotated as both RAC and REL to reflect its ethnoreligious nature (Litt, 1961). We also improve regulatory alignment by disambiguating GEN and SO (e.g., transgender bias labeled as GEN) and denoting pregnancy-targeted bias under GEN rather than PHY (Parrish et al., 2022). Labels outside our taxonomy (e.g., victim (Sap et al., 2020)) are excluded.
@@
The resulting dataset contains 46,781 entries, substantially larger than comparable benchmarks (e.g., 11,004 samples in (Wang et al., 2024)). Biased instances are more prevalent (around 70%), with most targeting a single demographic axis and roughly 12% of biased instances targeting multiple axes simultaneously. Among demographic targets, GEN, RAC, SO, and REL are most common, while PHY is least prevalent. Multi-axis biases most frequently combine {GEN, SO} or {GEN, RAC}.
@@ 3.3 Methodological testbed
3.3.1 Prompting
Brown et al. (2020) demonstrated that instruction-tuned LLMs can effectively perform a variety of tasks through textual prompting in zero-shot scenarios. Following (Inan et al., 2023; Palla et al., 2025), our evaluation framework employs policy-based prompting for bias detection. Specifically, the prompt includes a policy detailing the bias detection task and our demographic-based social bias taxonomy. To enhance the bias detection capabilities within this prompting framework, we also assess the benefits of incorporating few-shot examples for in-context learning. Specifically, we utilize the retrieval-augmented (RAG) framework (Chen et al., 2024a), where an embedding model identifies the most relevant in-context examples for each input instance from the training/development set.
Models. We consider several instruction-tuned models ranging from 8B to 72B parameters, e.g., GLM-4 (GLM et al., 2024), Llama-3.1 (Dubey et al., 2024), and Qwen-2.5 (Yang et al., 2024). We also analyze the guardrail model Llama Guard-3 (Inan et al., 2023) to explore if such models could directly be applied for general text bias detection. To perform RAG-based in-context sample selection, we use the BGE-M3 (Chen et al., 2024b) model.
3.3.2 Fine-tuning